OpenStudy (anonymous):

A sum of money is invested at 12% compounded quarterly. About how long will it take for the amount of money to double? Compound interest formula: V(T)=P(1+r/n)^nt t = years since initial deposit n = number of times compounded per year r = annual interest rate (as a decimal) P = initial (principal) investment V(t) = value of investment after t years p.S: Please explain, I have a couple problems like this so having an example to go off of would be nice. :) x Thanks

jimthompson5910 (jim_thompson5910):

Do you understand the formula and the variables for it?

OpenStudy (anonymous):

Kind of..

jimthompson5910 (jim_thompson5910):

If we want some sum of money (call it P dollars) to double, how much money do we want to end up with?

OpenStudy (anonymous):

Double the amount?

jimthompson5910 (jim_thompson5910):

yeah you start with P and it doubles to ???

OpenStudy (anonymous):

the 12%? ..

jimthompson5910 (jim_thompson5910):

that's the interest rate

jimthompson5910 (jim_thompson5910):

Let's say you start off with P = 100 dollars That doubles to $200 so what they want is the time it takes to go from $100 to $200

jimthompson5910 (jim_thompson5910):

to do that, you solve for t in \[\Large V(t) = P\left(1+\frac{r}{n}\right)^{n*t}\] \[\Large 200 = 100\left(1+\frac{0.12}{4}\right)^{4*t}\]

OpenStudy (anonymous):

But I'm confused because it doesn't give me sum to plug in, how am I supposed to know what to double?

jimthompson5910 (jim_thompson5910):

you can start off with any amount, it doesn't matter

jimthompson5910 (jim_thompson5910):

the doubling time does NOT change based on the starting amount the doubling time depends on the interest rate and the compounding frequency

OpenStudy (anonymous):

Thank you for trying to help, I think i'll just ask my teacher tomorrow though.

jimthompson5910 (jim_thompson5910):

alright, that works